7/Dec/99 Evening Shift

Peter Shanahan (shanahan@fnal.gov)
Tue, 7 Dec 1999 23:55:37 -0600

Shift Summary for eve shift, 7/Dec/99

On Shift: P. Shanahan

Hours of beam available at all intensities: 0.00
Hours of beam available at nominal intensity: 0.00
Hours of E832 data (special runs not included): 0.00
Hours of special runs (laser scans, muon runs...): 0.75

Accounting of lost time:
Beam - it doesn't work.

========================================================================
Runs ended this shift:
Run 15156 - LASER SCAN

Tapes brought to FCC this shift: None

=======================================================================
Accesses during shift: Open access. Work done on calorimeter to replace
HV cable for channel 2893.

========================================================================
Overview of shift: No beam during replacement of magnet at D4.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
16:01 PSH kills onsplit GUI, for more efficient web
browsing. ST consents.
16:15 JW returns HV dividers for 2892 and 2893 to correct
configuration after test.
Like a photon, JW repairs dynode connector for CsI
1136, at the same time.
18:10 DN puts 2 new Ar-Ethane bottles online on east side.
18:20 RR finished replacement of CsI 2893 HV cable.
15156 1 18:25 LASER SCAN!
19:40- Hourly silencing of VP marks torturous passage
23:59 of solitary confinement, without so much as a
hallucination of a visitor.
(Except MA and JS, checking up on special split).
23:30 PSH unplugs holiday lights for overnight safety.

========================================================================
Follow up Notes on Previous Shift

CsI 2446 has 2.54 count ped drift with respect to DB.

========================================================================
Items to Follow up on Next Shift

Watch CsI 2446